### v1

# Is the type-D NUT C-metric really “missing” from the most general Plebański-Demiański solution?

Abstract
It remains a long-standing problem, unsettled for almost two decades in General Relativity community, ever since Griffiths and Podolsky demonstrated in their previous paper [J.B. Griffiths and J. Podolsky, Classical Quantum Gravity 22, 3467 (2005)] that the type-D NUT C-metric seems to be absent from the most general family of the type-D Plebański-Demiański (P-D) solution. However, Astorino, in his recent article [M. Astorino, Phys. Rev. D 109, 084038 (2024)] presented a different form of rotating and accelerating black holes and showed that all known four-dimensional type-D accelerating black holes (without the NUT charge) can be recovered via various different limits in a definitive fashion. In particular, he provided, for the first time, the correct expressions for the type-D static accelerating black holes with a nonzero NUT charge, which was previously impossible using the traditional parametrization of the familiar P-D solution. Nevertheless, it still remains elusive that how these two different forms of the four-dimensional rotating and accelerating solutions are related. In this paper, we aim to fill this gap by finding the obvious coordinate transformations and parameter identifications between the vacuum metrics after two different parameterizations of the generated solution via the inverse scattering method from the seed metric – the Rindler vacuum background. We then resolve this “missing” puzzle by providing another Möbius transformation and linear combinations of the Killing coordinates, which clearly cast the type-D NUT C-metric into the familiar form of the P-D solution. Additionally, we propose an alternative new routine for the normalization of the obtained metric derived via the inverse scattering method from the vacuum seed solution, which could be potentially useful for the construction of higher-dimensional solutions using the trivial vacuum background as the seed metric.
I Introduction
The most general type-D solution in General Relativity has been well-known for more than half a century JMP10-1195 ; AP90-196 ; AP98-98 , and is often referred to as the Plebański-Demiański (P-D) metric. With the advent of the conveniently factorized form CQG22-109 of two quartic structure functions, this solution has subsequently been extensively studied during the past twenty years, see the book GP2009 for a comprehensive review. However, as demonstrated in Ref. CQG22-3467 , the type-D NUT C-metric is apparently accelerating and can be transformed into the usual Taub-NUT solution. Since then, “such a solution has to date neither been identified nor proved not to exist.” This raised a “missing” puzzle for almost two decades, as the type-D NUT C-metric appears to be absent from the most general family of the type-D P-D solution.
Recently, Astorino PRD109-084038 presented a different form of the rotating and accelerating black holes that belong to the type-D family (although no obvious proof of this issue is provided in that paper) and showed that all known four-dimensional type-D accelerating black holes without the NUT charge can be obtained via various different limits in a definitive fashion. Moreover, he presented, for the first time, the correct expressions for the type-D NUT C-metric, which are not available in the traditional parametrization for the familiar P-D solution. Lately, this solution has also been generalized in 2404.06551 to include the electro-magnetic charges and a nonzero cosmological constant.
As far as the NUT C-metric is concerned, apart from the above type-D one, there is also another kind of the accelerating Taub-NUT solution PRD74-084031 , which belongs to a more general type-I PRD102-084024 that has attracted some recent interest PRD108-124025 ; PRD108-024059 ; 2309.13656 . Although the initial form of this solution is complicated and has been obtained via a hybrid solution-generating method from a specific seed, in fact it can be simply generated PRD108-124025 ; PRD108-024059 via the Ehler transformation from the C-metric. The Ehler transformation not only endows the black hole horizon with the NUT charge, but also allocates a port of the NUT charge to the accelerating horizon. From the perspective of the double Kerr-Schild formalism, the type-D NUT C-metric can be viewed as a perturbation on the pure Rindler background. It seems that one could regard the NUT-charged Rindler metric as the background spacetime of the type-I NUT C-metric. However, it is unclear whether the type-I C-metric admits a similar double Kerr-Schild formulation. Based on the same reasoning, when the Harrison transformation is applied to the seed C-metric, one gets a type-I electrically charged C-metric, whose background is the electrified Rindler vacuum. However, a well-known type-D charged static accelerating solution (Reissner-Nordström C-metric, or RN C-metric) is already included in the most general P-D family. Like the Ehler transformation that allocates a port of the NUT charge to the Rindler background, the Harrison transformation also assigns a port of the electric charge to the Rindler background. Thus, in the type-I RN C-metric, the accelerating horizon acquires the electric charge; in other words, its background is a charged Rindler metric. In this sense, the original Kinnersley-Walker’s interpretation PRD2-1359 of the uniformly accelerating charged mass as two “oppositely” accelerated black holes apart from each other seems not to be very accurate, as it can not distinguish the type-D accelerating black holes from the type-I ones. Rather, one could say that the type-D accelerating solution is a black hole siting on the pure Rindler background, while the type-I accelerating metric is that on the more general Rindler background with dynamical parameters (the NUT charge, electro-magnetic charge) plus other kinetic parameters (such as the rotation parameter and possibly the cosmological constant). It is noted that while the type-D solution admits an extension to include a nonzero cosmological constant, the type-I metric seems unable to permit such a generalization. By the way, we mention that this is also another kind of the NUT C-metric presented in Ref. PRD91-124005 , which is dubbed the [4-1]-soliton solution PRD93-044021 . This solution still remains unexplored to date, but we can anticipate it to be of type-I in the Petrov classification.
What remains opaque in the Astorino’s work is how these two different forms of the four-dimensional rotating and accelerating solutions are related, and why the type-D NUT C-metric is “missing” from the P-D metric. The main purpose of this paper is to resolve this “missing” puzzle in two different ways. First, after generating the rotating and accelerating Kerr-NUT metric via the inverse scattering method (ISM) JETP48-985 ; JETP50-1 ; BV2005 from the seed Rindler vacuum background, we obtain two distinct forms of the same solution by using two different parameterizations. This clearly shows that both metrics are isometric to each other. Then, we find the obvious Möbius transformations and parameter identifications between them. In the meanwhile, we will illustrate what limit should be taken in the general P-D solution to obtain the type-D NUT C-metric. In particular, we shall present another Möbius transformation and linear combinations of the Killing coordinates that clearly cast the type-D NUT C-metric into the familiar form of the P-D solution, with the quartic structure function being factorized into two quadratic functions in a manner different from that of Griffiths-Podolsky’s.
The main parts of our paper is organized as follows. In Sec. II, we present the ISM construction of the type-D accelerating Kerr-NUT metric as a 2-soliton transformation from the Rindler vacuum background, and two distinct parameterizations to arrive at two different rational expressions of the same solution. As preparation, we take the C-metric as a simple example to compare the double Kerr-Schild formalism with the ISM perspective, and attempt to combine these two seemly unrelated schemes. This idea has never been advocated in the past years. Then after a minor retrospect of the ISM algorithm, we devise a new strategy for the normalization of the obtained metric derived via the ISM construction from the vacuum seed solution, which could facilitate the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. Our normalization scheme is corroborated in the Appendix Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme by deriving the 5-dimensional Schwarzschild-Tangherlini solution NC27-636 from the flat vacuum background in two different ways. The remaining two sections III and IV are entirely new and constitute the main contribution of this work, where Möbious transformations of the radial and angular coordinates, and linear combinations of the Killing coordinates as well as various parameter identifications between these two metrics are plainly displayed. Our paper is ended up with a brief summary of this work, followed by the appendix with two simple examples to illuminate our new normalization strategy.
II Accelerating Kerr-NUT solution: ISM construction and two different parameterizations
II.1 The C-metric and its seed
We begin with two different descriptions of the C-metric, whose line element in spherical-like coordinates ($t,r,x=\cos\theta,\bar{\phi}$)
$d\tilde{s}^{2}=\frac{1}{(1+\alpha\,rx)^{2}}\Big{[}-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\frac{dx^{2}}{h(x)}+r^{2}h(x)d\tilde{\phi}^{2}\Big{]}\,,$ (1)
can be rewritten in terms of the soliton formalism as
$d\tilde{s}^{2}\equiv\alpha\,ds^{2}=-\frac{\mu_{1}\mu_{3}}{\mu_{2}}dt^{2}+\frac{\rho^{2}\mu_{2}}{\mu_{1}\mu_{3}}d\phi^{2}+\frac{\mu_{1}\mu_{3}R_{12}^{2}R_{23}^{2}}{(1-2m\alpha)^{2}\mu_{2}R_{13}^{2}R_{11}R_{22}R_{33}}\big{(}d\rho^{2}+dz^{2}\big{)}\,,$ (2)
where the Weyl coordinates are defined as
$\rho=\frac{\sqrt{r^{2}f(r)h(x)}}{(1+\alpha\,rx)^{2}}\,,\quad z=\frac{(\alpha\,r+x)[(1+m\alpha\,x)r-m]}{(1+\alpha\,rx)^{2}}\,,$
in which two structure functions being
$f(r)=\big{(}1-\alpha^{2}r^{2}\big{)}\Big{(}1-\frac{2m}{r}\Big{)}\,,\qquad h(x)=\big{(}1-x^{2}\big{)}(1+2m\alpha\,x)\,.$ (3)
In the above, we have also made a rescaling $\tilde{\phi}=\alpha\phi$ and introduced the following notations:
$R_{kl}=\rho^{2}+\mu_{k}\mu_{l}\,,\quad\mu_{k}=\sqrt{\rho^{2}+(z-z_{k})^{2}}-(z-z_{k})\,,$ (4)
where the 3-soliton $\mu_{k}$’s are algebraically expressed as
$\mu_{1}=\frac{(1-\alpha\,r)(r-2m)(1-x)}{(1+\alpha\,rx)^{2}}\,,\quad\mu_{2}=\frac{(1-\alpha\,r)r(1-x)(1+2m\alpha\,x)}{(1+\alpha\,rx)^{2}}\,,\quad\mu_{3}=\frac{(1-\alpha^{2}r^{2})(1+2m\alpha\,x)}{\alpha(1+\alpha\,rx)^{2}}\,,$ (5)
which are associated with three rod endpoints: $z_{1}=-m\,,z_{2}=m\,,z_{3}=1/(2\alpha)$, respectively.
The zero-mass limit $\mu_{2}\mapsto\mu_{1}$ is the Rindler vacuum
$d\bar{s}^{2}\equiv\alpha\,ds^{2}=-\mu_{3}dt^{2}+\frac{\rho^{2}}{\mu_{3}}d\phi^{2}+\frac{\mu_{3}}{\rho^{2}+\mu_{3}^{2}}\big{(}d\rho^{2}+dz^{2}\big{)}\,,$ (6)
which will be used later as our seed metric. The Rindler metric can be viewed as 1-soliton transformation from the flat Minkowski background with the metric components: $g_{\mu\nu}=(-1,\rho^{2},1,1)$. However, it is noted that the Rindler vacuum is trivial and is entirely equivalent to flat Minkowski metric since they are related via some simple coordinate transformations.
So in the soliton formalism, the C-metric can be interpreted either as a 3-soliton transformation to the flat Minkowski background or a 2-soliton transformation to the Rindler background. In this article, we will take the latter viewpoint and derive the accelerating Kerr-NUT solution via the ISM procedure by using the Rindler background (6) as the seed metric. Note that usually the seed solution need not to be exactly the background metric and frequently can be taken as more complicated even singular one for the technical aim. However, a simple background metric is surely more preferable in practice as a seed solution.
On the other hand, the C-metric (1) admits the double Kerr-Schild formulation as follows:
$d\tilde{s}^{2}$ $=$ $\frac{1}{(1+\alpha\,rx)^{2}}\Big{[}-\big{(}1-\alpha^{2}r^{2}\big{)}d\bar{t}^{2}+\frac{dr^{2}}{1-\alpha^{2}r^{2}}+r^{2}\frac{dx^{2}}{1-x^{2}}+r^{2}\big{(}1-x^{2}\big{)}d\bar{\phi}^{2}\Big{]}$ (7)
$+\frac{2m\big{(}1-\alpha^{2}r^{2}\big{)}}{r(1+\alpha\,rx)^{2}}\Big{(}d\bar{t}+\frac{dr}{1-\alpha^{2}r^{2}}\Big{)}^{2}+\frac{2m\alpha\,x\big{(}1-x^{2}\big{)}r^{2}}{(1+\alpha\,rx)^{2}}\Big{(}d\bar{\phi}+\frac{I\,dx}{1-x^{2}}\Big{)}^{2}\,.$
after making the coordinate transformations
[…]
The zero-mass part is just the Rindler vacuum written in the C-metric coordinates ($x,y$) with $y=-1/(\alpha\,r)$. In view of this, the C-metric can be interpreted as a perturbation around the Rindler background. Incidentally, it should be pointed out that the above Kerr-Schild ansatz is completely applicable to the most general P-D metric AP98-98 .
However, it seems that the following double quasi-Kerr-Schild prescription
$d\tilde{s}^{2}$ $=$ $\frac{1}{(1+\alpha\,rx)^{2}}\Big{[}-\big{(}1-\alpha^{2}r^{2}\big{)}dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}\frac{dx^{2}}{h(x)}+r^{2}\big{(}1-x^{2}\big{)}d\tilde{\phi}^{2}\Big{]}$ (8)
$+\frac{2m\big{(}1-\alpha^{2}r^{2}\big{)}}{r(1+\alpha\,rx)^{2}}dt^{2}+\frac{2m\alpha\,x\big{(}1-x^{2}\big{)}r^{2}}{(1+\alpha\,rx)^{2}}d\tilde{\phi}^{2}\,,$
is in more accordance with the above 2-soliton viewpoint by virtue of the 2-dimensional metrics are conformal to each other:
$\frac{1}{(1+\alpha\,rx)^{2}}\Big{[}\frac{dr^{2}}{f(r)}+r^{2}\frac{dx^{2}}{h(x)}\Big{]}\simeq C_{f}(\rho,z)(d\rho^{2}+dz^{2})\,.$
II.2 ISM: A minor preliminary
In a $D$-dimensional space-time with $(D-2)$ Killing vectors, if its metric, whose components depend on the Weyl coordinates $\rho$ and $z$ only, is written as PRD70-124002
$ds^{2}=g_{ab}dx^{a}dx^{b}+f(\rho,z)(d\rho^{2}+dz^{2})\,,$ (9)
and supplemented with the canonical condition:
$\det\mathbf{g}=-\rho^{2}\,,$ (10)
then the vacuum Einstein equations are equivalent to the following equations for the $(D-2)\times(D-2)$ matrix $\mathbf{g}=(g_{ab})$:
$\partial_{\rho}\mathbf{U}+\partial_{z}\mathbf{V}=0\,,$ (11)
and for the conformal factor $f(\rho,z)$, which can be integrated by quadratures:
$\begin{split}\partial_{\rho}\ln\,f&=-\frac{1}{\rho}+\frac{1}{4\rho}\mathop{\rm Tr}\nolimits(\mathbf{U}^{2}-\mathbf{V}^{2})\,,\\ \partial_{z}\ln\,f&=\frac{1}{2\rho}\mathop{\rm Tr}\nolimits(\mathbf{U}\mathbf{V})\,,\end{split}$ (12)
where the following two matrices are introduced
$\mathbf{U}=\sqrt{-g}(\partial_{\rho}\mathbf{g})\mathbf{g}^{-1}\,,\qquad\mathbf{V}=\sqrt{-g}(\partial_{z}\mathbf{g})\mathbf{g}^{-1}\,.$ (13)
System (11) is completely integrable and can be viewed as the compatibility condition of the Lax pair for the $(D-2)\times(D-2)$ generating matrix $\mathbf{\Psi}(\lambda,\rho,z)$:
$\begin{split}&\partial_{\rho}\mathbf{\Psi}+\frac{2\lambda\rho}{\lambda^{2}+\rho^{2}}\partial_{\lambda}\mathbf{\Psi}=\frac{\rho\mathbf{U}+\lambda\mathbf{V}}{\lambda^{2}+\rho^{2}}\mathbf{\Psi}\,,\\ &\partial_{z}\mathbf{\Psi}-\frac{2\lambda^{2}}{\lambda^{2}+\rho^{2}}\partial_{\lambda}\mathbf{\Psi}=\frac{\rho\mathbf{V}-\lambda\mathbf{U}}{\lambda^{2}+\rho^{2}}\mathbf{\Psi}\,,\end{split}$ (14)
with $\lambda$ being a complex spectral parameter independent of $\rho$ and $z$. When $\lambda=0$, one gets $\mathbf{g}(\rho,z)=\mathbf{\Psi}(0,\rho,z)$. One can construct new solitonic solutions from a seed metric via the dressing procedure:
$\mathbf{\Psi}=\chi\!\!\!\!\chi\,\mathbf{\Psi}_{0}\,,\qquad\chi\!\!\!\!\chi=\mathbf{I}+\sum_{k=1}^{n}\frac{\mathbf{n}^{{\dagger}(k)}\mathbf{m}^{(k)}}{\lambda-\mu_{k}}\,,$ (15)
in which $\mathbf{n}^{{\dagger}(k)}$ denotes the transpose of the BZ vector $\mathbf{n}^{(k)}$ defined below, $\mu_{k}$ represents a soliton with its corresponding anti-soliton being $\nu_{k}=-\rho^{2}/\mu_{k}$, and the constant $z_{k}$ is the turning point (or rod endpoint) on the $z$ axes.
The expression for the generated matrix of the general $n$-soliton solutions is:
$\widetilde{\mathbf{g}}=\mathbf{g}_{0}-\sum_{k,l=1}^{n}\frac{\mathbf{l}^{{\dagger}(k)}(\mathbf{\Gamma}^{-1})_{kl}\mathbf{l}^{(l)}}{\mu_{k}\mu_{l}}\,.$ (16)
with the following notations being introduced
$\mathbf{\Gamma}_{kl}=\frac{\mathbf{m}^{(k)}\mathbf{g}_{0}\mathbf{m}^{{\dagger}(l)}}{\rho^{2}+\mu_{k}\mu_{l}}\,,\qquad\mathbf{l}^{(k)}=\mathbf{m}^{(k)}\mathbf{g}_{0}\,,\qquad\mathbf{m}^{(k)}=\mathbf{m}_{0}^{(k)}\mathbf{\Psi}_{0}^{-1}(\mu_{k},\rho,z)\,,\qquad\mathbf{n}^{(k)}=\sum_{l=1}^{n}\frac{(\mathbf{\Gamma}^{-1})_{kl}\mathbf{l}^{(l)}}{\mu_{l}}\,,$ (17)
where the BZ vectors $\mathbf{m}_{0}^{(i)}$ consist of arbitrary constants with one of which can be set to unity for each vector.
In the case of a diagonal seed, the generating matrix $\mathbf{\Psi}_{0}(\mu_{i},\rho,z)$ can be easily obtained from the seed metric $\mathbf{g}_{0}(\rho,z)$ by simply replacing every soliton or anti-soliton via $\mu_{k}\to\mu_{k}-\lambda$ and $\rho^{2}\to\rho^{2}-2\lambda\,z-\lambda^{2}$, or $\rho^{2}/\mu_{k}\to\rho^{2}/\mu_{k}+\lambda$.
In general, the generated metric $\widetilde{\mathbf{g}}$ is unphysics due to the fact that it does not fulfill the canonical condition (10): $\det\tilde{\mathbf{g}}\not=-\rho^{2}$, so it should be appropriately normalized.
As far as the normalization scheme is concerned, to the best of our knowledge, there exist three different kinds of strategies up to date. (I) Uniform normalization BV2005 : Like the four-dimensional case, all the metric components of the Killing parts in the higher-dimensional case are multiplied by the same normalization factor. The main defect of this method lies in that it yields some fractional singularities on the symmetry axes, so that the final solution has little physics meaning. (II) Partial block uniform normalization PRD73-064009 ; PRD73-124034 ; PRD74-104004 ; PTP116-319 , which stems from early work PRD34-2990 ; PLA136-269 : This method only applies the above uniform normalization to a part of the block diagonal matrix of the generated metric, so it is only effective in the case when not all rotation parameters are turned on, for instance, it is limited to the singly rotating case in five dimensions, and so on. (III) Pomeransky’s remove–re-add tricky PRD73-044004 : If the seed metric already satisfies the canonical condition, then after removing and readding the same number of solitions or anti-solitons, the generated metric automatically subjects to the canonical condition also. At the same time, the conformal factor can be simply computed as $f=f_{0}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}_{0}}$, where $\mathbf{\Gamma}_{0}$ can be obtained from $\mathbf{\Gamma}$ by setting the nontrivial BZ constants to zero. This is a wise strategy that is widely adopted nowadays, for example, the doubly rotating Myers-Perry solution AP172-304 is re-generated via this method from the five-dimensional Schwarzschild-Tangherlini solution NC27-636 . However, there is a fly in the ointment in that it can not work effective in the case when the seed solution is a vacuum background metric where there is no soliton that can be removed. In particular, it is claimed PRD73-044004 that the five-dimensional Schwarzschild-Tangherlini solution NC27-636 cannot be obtained as a 2-soliton solution on the flat Minkowski background.
Below we propose the fourth scheme for the normalization of the obtained metric derived from the vacuum seed solution, which would be potentially useful for the construction of higher dimensional solutions by using the trivial vacuum background as the seed metric. In the appendix, we will present two simple examples to illuminate this method. First, let’s consider a simple ansatz:
$\mathbf{g}=\mathbf{W}\widetilde{\mathbf{g}}\mathbf{W}\,,\qquad\mathbf{W}=\mathop{\rm diag}\nolimits(w_{1}\,,\cdots\,,w_{D-2})\,,$ (18)
by simply employing a diagonal matrix $\mathbf{W}$ with which each of its entry satisfies the Laplace equation:
$\nabla^{2}\ln\,w_{i}=\Big{(}\partial_{\rho}^{2}+\frac{1}{\rho}\partial_{\rho}+\partial_{z}^{2}\Big{)}\ln\,w_{i}=0\,.$ (19)
The solution of each entry in the above equation has the general form: $w_{i}=\rho^{b_{i}}\mu_{i}^{c_{i}}$, in which $b_{i}$ and $c_{i}$ are arbitrary constants. The entries in $\mathbf{W}$ are chosen so that $-\rho^{2}=w_{1}^{2}\cdots\,w_{D-2}^{2}\det\widetilde{\mathbf{g}}$ is satisfied. By virtue of the fact that $\det\widetilde{\mathbf{g}}$ is proportional to $\det\mathbf{g}_{0}$, actually one only needs to choose different powers in $w_{i}$.
If the generated metric is still diagonal just as its seed metric, then the above means amounts to $\mathbf{g}=\mathbf{W}^{2}\widetilde{\mathbf{g}}$, and it is easy to compute the conformal factor. On the other hand, if all the diagonal entries are identical $w_{i}=w$, then the above scheme is equivalent to the original uniform normalization strategy. Supposed that only some of the diagonal entries are set to be identical, then this amounts to the partial block uniform normalization. Finally, if all $w_{i}$’s are appropriately chosen such as ($-\rho^{2}/\mu_{i}^{2}$) or ($-\mu_{i}^{2}/\rho^{2}$), this scheme is actually equivalent to removing solition or anti-soliton at least in the case of the generated metric being static (diagonal).
Second, let’s handle the more general case. Since, the generated metric in general still needs to make further linear combinations of the Killing coordinates, that is, $\hat{\mathbf{g}}=\mathbf{A}^{{\dagger}}\mathbf{g}\mathbf{A}$, where the rotation matrix $\mathbf{A}$ is unimodular and its every entry is a constant. Then, we can have $\hat{\mathbf{g}}=\hat{\mathbf{W}}^{{\dagger}}\widetilde{\mathbf{g}}\hat{\mathbf{W}}$, where $\hat{\mathbf{W}}=\mathbf{W}\mathbf{A}$, $\hat{\mathbf{W}}^{{\dagger}}$ is the matrix transpose of $\hat{\mathbf{W}}$ and each of its entry obeys the above Laplace equation (19). Clearly it is a trouble thing to choose a suitable matrix $\hat{\mathbf{W}}$ so that the final metric is the expected one. However, one need not to worry about this, since the Pomeransky’s remove–re-add tricky is already very effective to generate the rotating solution, so our method just plays a nice complementary where it fails to work.
II.3 ISM construction of the accelerating Kerr-NUT solution
Although the accelerating Kerr-NUT solution can be derived as a 3-soliton transformation via the ISM procedure from the flat Minkowski background, in accordance with the above double Kerr-Schild perspective, we prefer to rederive it as a 2-soliton transformation by using the Rindler background (6) as the seed metric, whose Killing part is
$\mathbf{g}_{0}=\mathop{\rm diag}\nolimits(-\mu_{3}\,,\rho^{2}/\mu_{3})=\mathop{\rm diag}\nolimits(-\mu_{3}\,,-\nu_{3})\,,$ (20)
with the conformal factor $f_{0}=\mu_{3}/R_{33}$, so that the generating matrix is easily obtained as
$\mathbf{\Psi}_{0}=\mathop{\rm diag}\nolimits(\lambda-\mu_{3}\,,\lambda-\nu_{3})=\mathop{\rm diag}\nolimits(\lambda-\mu_{3}\,,\lambda+\rho^{2}/\mu_{3})\,.$ (21)
Now we perform a 2-soliton transformation to this Rindler background: add a soliton $\mu_{1}$ at $z=z_{1}$ with the vector $m_{0}^{(1)}=(C_{1}\,,1)$ and a soliton $\mu_{2}$ at $z=z_{2}$ with the vector $m_{0}^{(2)}=(1\,,C_{2})$, respectively. Note that we take the 2-soliton in the order: $\mu=(\mu_{1}\,,\mu_{2})$, so that the generated metric exactly reproduces the same form (2) for the C-metric when the BZ parameters are set to zero: $C_{1}=C_{2}=0$. Then two BZ vectors constitute a 2× $\times\,2$ 2 matrix:
$\mathbf{m}=\bigg{(}\begin{array}[]{cc}\frac{C_{1}}{\mu_{13}}&\frac{\mu_{3}}{R_{13}}\\ \frac{1}{\mu_{23}}&\frac{C_{2}\mu_{3}}{R_{23}}\end{array}\bigg{)}\,,\quad\mathbf{l}=\bigg{(}\begin{array}[]{cc}\frac{-C_{1}\mu_{3}}{\mu_{13}}&\frac{\rho^{2}}{R_{13}}\\ \frac{-\mu_{3}}{\mu_{23}}&\frac{C_{2}\rho^{2}}{R_{23}}\end{array}\bigg{)}$ (22)
where a short notation is included for briefness: $\mu_{ij}=\mu_{i}-\mu_{j}$. The $\mathbf{\Gamma}$ matrix is also recorded as follows:
$\mathbf{\Gamma}=\bigg{(}\begin{array}[]{cc}\frac{-C_{1}^{2}\mu_{3}}{\mu_{13}^{2}R_{11}}+\frac{\mu_{3}\rho^{2}}{R_{13}^{2}R_{11}}&\frac{-C_{1}\mu_{3}}{\mu_{13}\mu_{23}R_{12}}+\frac{C_{2}\mu_{3}\rho^{2}}{R_{12}R_{13}R_{23}}\\ \frac{-C_{1}\mu_{3}}{\mu_{13}\mu_{23}R_{12}}+\frac{C_{2}\mu_{3}\rho^{2}}{R_{12}R_{13}R_{23}}&\frac{-\mu_{3}}{\mu_{23}^{2}R_{22}}+\frac{C_{2}^{2}\mu_{3}\rho^{2}}{R_{23}^{2}R_{22}}\end{array}\bigg{)}\,.$ (23)
The final physics metric can be obtained via either the uniform normalization scheme or our prescription proposed in the above
$\mathbf{g}=\mathbf{W}\widetilde{\mathbf{g}}\mathbf{W}=-\frac{\mu_{1}\mu_{2}}{\rho^{2}}\widetilde{\mathbf{g}}\,,$ (24)
in which we have taken $\mathbf{W}=\mathop{\rm diag}\nolimits(\sqrt{-\mu_{1}\mu_{2}}/\rho\,,\sqrt{-\mu_{1}\mu_{2}}/\rho)$. The conformal factor is
$f=-16k^{2}f_{0}\frac{\mu_{1}^{3}\mu_{2}^{3}}{\rho^{2}\mu_{12}^{2}}\det\mathbf{\Gamma}=16k^{2}\frac{B}{H}\,.$ (25)
The generated metric represents the accelerating Kerr-NUT solution, whose line element reads
$ds^{2}=\frac{-\mu_{3}A}{\mu_{1}\mu_{2}B}\Big{(}dt+\frac{\mu_{12}C}{\mu_{3}A}d\phi\Big{)}^{2}+\frac{\rho^{2}\mu_{1}\mu_{2}B}{\mu_{3}A}d\phi^{2}+16k^{2}\frac{B}{H}\big{(}d\rho^{2}+dz^{2}\big{)}\,,$ (26)
where
$A$ $=$ $\big{[}C_{1}(C_{2}\mu_{2}\mu_{23}R_{12}-\rho\mu_{12}R_{23})R_{13}-\mu_{1}\mu_{13}(C_{2}\rho\mu_{2}\mu_{12}\mu_{23}+R_{12}R_{23})\big{]}$
×$\times\big{[}C_{1}(C_{2}\mu_{2}\mu_{23}R_{12}+\rho\mu_{12}R_{23})R_{13}+\mu_{1}\mu_{13}(C_{2}\rho\mu_{2}\mu_{12}\mu_{23}-R_{12}R_{23})\big{]}\,,$
$B$ $=$ $C_{1}^{2}\big{(}C_{2}^{2}\mu_{23}^{2}R_{12}^{2}+\mu_{12}^{2}R_{23}^{2}\big{)}R_{13}^{2}-2C_{1}C_{2}\mu_{13}\mu_{23}R_{13}R_{23}R_{11}R_{22}$
$+\big{(}C_{2}^{2}\rho^{4}\mu_{12}^{2}\mu_{23}^{2}+R_{12}^{2}R_{23}^{2}\big{)}\mu_{13}^{2}\,,$
$C$ $=$ $C_{1}\mu_{2}\mu_{13}\big{(}C_{2}^{2}\rho^{2}\mu_{23}-R_{23}^{2}\big{)}R_{11}R_{12}R_{13}+C_{2}\mu_{1}\mu_{23}\big{(}C_{1}^{2}R_{13}^{2}-\rho^{2}\mu_{13}^{2}\big{)}R_{12}R_{22}R_{23}\,,$
$H$ $=$ $\frac{\mu_{12}^{2}\mu_{13}^{2}\mu_{23}^{2}}{\mu_{1}^{3}\mu_{2}^{3}\mu_{3}^{3}}R_{12}^{2}R_{13}^{2}R_{23}^{2}R_{11}R_{22}R_{33}\,.$
When $C_{1}=C_{2}=0$, the above generated 2-soliton solution just reduces to the C-metric (2).
Note that the above accelerating Kerr-NUT solution can also be regenerated via the “remove–re-add” tricky from the C-metric (2). The process is briefly sketched as follows. We first remove an anti-soliton $\nu_{1}$ at $z=z_{1}$ and a soliton $\mu_{2}$ at $z=z_{2}$ both with the trivial vectors $(1,0)$, then rescale the metric by a factor $\mu_{1}/\mu_{2}$. This essentially generates the above Rindler seed (20):
$\mathbf{g}_{0}=\frac{\mu_{1}}{\mu_{2}}\mathop{\rm diag}\nolimits\Big{(}\frac{\mu_{2}^{2}}{\mu_{1}^{2}}\,,1\Big{)}\times\mathop{\rm diag}\nolimits\Big{(}-\frac{\mu_{1}\mu_{3}}{\mu_{2}}\,,\frac{\rho^{2}\mu_{2}}{\mu_{1}\mu_{3}}\Big{)}\,.$
Next we re-add the anti-soliton $\nu_{1}$ with the vector $m_{0}^{(1)}=(1,-C_{1}$) and the soliton $\mu_{2}$ with the vector $m_{0}^{(2)}=(1,C_{2}$), and multiple the generated metric by a factor $\mu_{2}/\mu_{1}$ so that the final metric is physics because it satisfies the canonical condition: $\det\mathbf{g}=-\rho^{2}$. The conformal factor can be easily evaluated via $f=f_{00}\det\mathbf{\Gamma}/{\det\mathbf{\Gamma}(C_{1}=0,C_{2}=0)}$, where $f_{00}$ is the conformal factor given by Eq. (2).
II.4 Astorino’s parametrization
We first transform from Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to spherical-like coordinates ($r,x$) via:
$\rho=\frac{\sqrt{\Delta(r)P(x)}}{(1+\alpha\,rx)^{2}}\,,\qquad z=\frac{(x+\alpha\,r)\big{[}(r-m)(1+m\alpha\,x)+\sigma^{2}\alpha\,x\big{]}}{(1+\alpha\,rx)^{2}}\,,$
where
$\Delta(r)=\big{[}(r-m)^{2}-\sigma^{2}\big{]}\big{(}1-\alpha^{2}r^{2}\big{)}\,,\quad P(x)=\big{(}1-x^{2}\big{)}\big{[}(1+m\alpha\,x)^{2}-\sigma^{2}\alpha^{2}x^{2}\big{]}\,,$ (27)
in which $\sigma=\sqrt{m^{2}+n^{2}-a^{2}}$.
The locations of three turning points are taken to be
$z_{1}=-\sigma,\quad z_{2}=\sigma,\quad z_{3}=\frac{1-\alpha^{2}(m^{2}-\sigma^{2})}{2\alpha}\,,$
so their corresponding solitons are purely algebraic:
$\mu_{1}$ $=$ $\frac{(1-\alpha\,r)(r-m-\sigma)[1+(m-\sigma)\alpha\,x)](1-x)}{(1+\alpha\,rx)^{2}}\,,$
$\mu_{2}$ $=$ $\frac{(1-\alpha\,r)(r-m+\sigma)[1+(m+\sigma)\alpha\,x)](1-x)}{(1+\alpha\,rx)^{2}}\,,$ (28)
$\mu_{3}$ $=$ $\frac{\big{(}1-\alpha^{2}r^{2}\big{)}[1+(m-\sigma)\alpha\,x)][1+(m+\sigma)\alpha\,x)]}{\alpha(1+\alpha\,rx)^{2}}\,,$
which reduce to Eq. (5) exactly when the static C-metric limit is reached ($\sigma=m$).
It is very tedious to convert the four functions ($A,B,C,H$) into the expressions in terms of spherical-like coordinates. However, if we take
$C_{1}=\frac{(m-\sigma)[1+(m+\sigma)\alpha]}{(a-n)[1-(m-\sigma)\alpha]}\,,\quad C_{2}=\frac{(m-\sigma)[1-(m+\sigma)\alpha]}{(a+n)[1+(m-\sigma)\alpha]}\,,$ (29)
and
$k=\frac{(m+\sigma)[1+(m-\sigma)\alpha][1-(m-\sigma)\alpha]}{2\alpha^{3/2}\sqrt{1+\alpha^{2}a^{2}}}\,,$ (30)
then we are able to first write the 2-dimensional metric as
$f\big{(}d\rho^{2}+dz^{2}\big{)}=\frac{\Sigma}{\big{(}1+\alpha^{2}a^{2}\big{)}(1+\alpha\,rx)^{2}}\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}\,,$
in which
$\Sigma=r^{2}+(ax-n)^{2}+2n\alpha\,r\big{[}a\big{(}1+x^{2}\big{)}-2nx\big{]}+\alpha^{2}r^{2}(a-nx)^{2}+\big{(}a^{2}-n^{2}\big{)}^{2}\alpha^{2}x^{2}\,.$ (31)
Finally, after some cumbersome algebraic simplifications, the Killing parts can be converted into
$g_{ab}dx^{a}dx^{b}=\frac{-F}{(1+\alpha\,rx)^{2}\Sigma\alpha}\Big{(}dt-\frac{\omega\alpha}{F}d\varphi\Big{)}^{2}+\frac{\Sigma\Delta(r)P(x)\alpha}{(1+\alpha\,rx)^{2}F}d\varphi^{2}\,,$
where
$F$ $=$ $\big{[}1-\big{(}a^{2}-n^{2}\big{)}\alpha^{2}x^{2}\big{]}^{2}\Delta(r)-\big{[}a\big{(}1+\alpha^{2}r^{2}\big{)}+2n\alpha\,r\big{]}^{2}P(x)\,,$
$\omega$ $=$ $\big{[}a\big{(}1+x^{2}\big{)}-2nx\big{]}\big{[}1-\big{(}a^{2}-n^{2}\big{)}\alpha^{2}x^{2}\big{]}\Delta(r)+\big{(}r^{2}+n^{2}-a^{2}\big{)}\big{[}a\big{(}1+\alpha^{2}r^{2}\big{)}+2n\alpha\,r\big{]}P(x)\,.$
After rescaling the Killing coordinates as $t\mapsto\sqrt{\alpha}t\,,\varphi\mapsto\varphi/\sqrt{\alpha}$, the above two parts are collectively written in terms of Lewis-Papapetrou coordinates as PRD109-084038
$ds^{2}=\frac{-F}{(1+\alpha\,rx)^{2}\Sigma}\Big{(}dt-\frac{\omega}{F}d\varphi\Big{)}^{2}+\frac{\Sigma}{(1+\alpha\,rx)^{2}F}\bigg{\{}\Delta(r)P(x)d\varphi^{2}+\frac{F}{1+\alpha^{2}a^{2}}\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}\bigg{\}}\,,$ (32)
which can also be recast into Boyer-Lindquist-like form
$ds^{2}=\frac{1}{(1+\alpha\,rx)^{2}}\bigg{\{}\frac{-\Delta(r)X^{2}+P(x)Y^{2}}{\Sigma}+\frac{\Sigma}{1+\alpha^{2}a^{2}}\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}\bigg{\}}\,,$ (33)
where
$\begin{split}X&=\big{[}1-\big{(}a^{2}-n^{2}\big{)}\alpha^{2}x^{2}\big{]}dt+\big{[}2nx-a\big{(}1+x^{2}\big{)}\big{]}d\varphi\,,\\ Y&=\big{[}a\big{(}1+\alpha^{2}r^{2}\big{)}+2n\alpha\,r\big{]}dt+\big{(}r^{2}+n^{2}-a^{2}\big{)}d\varphi\,.\end{split}$ (34)
In the absence of the NUT charge, the above metric simply recovers the accelerating Kerr solution in the familiar form after further shifting the Killing coordinates:
$t\mapsto\frac{t-2a\varphi}{\sqrt{1+\alpha^{2}a^{2}}}\,,\quad\varphi\mapsto-\frac{\alpha^{2}at+\big{(}1-\alpha^{2}a^{2}\big{)}\varphi}{\sqrt{1+\alpha^{2}a^{2}}}\,.$
On the other hand, the accelerating NUT C-metric takes a simple expression:
$ds^{2}=\frac{1}{(1+\alpha\,rx)^{2}}\bigg{\{}-\frac{\Delta(r)}{\Sigma}\big{[}\big{(}1+n^{2}\alpha^{2}x^{2}\big{)}dt+2nxd\varphi\big{]}^{2}+\frac{P(x)}{\Sigma}\big{[}2n\alpha\,rdt+\big{(}r^{2}+n^{2}\big{)}d\varphi\big{]}^{2}+\Sigma\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}\bigg{\}}\,,$ (35)
in which
$\begin{split}\Delta(r)&=\big{(}r^{2}-2mr-n^{2}\big{)}\big{(}1-\alpha^{2}r^{2}\big{)}\,,\quad P(x)=\big{(}1-x^{2}\big{)}\big{(}1+2m\alpha\,x-n^{2}\alpha^{2}x^{2}\big{)}\,,\\ \Sigma&=\big{(}r^{2}+n^{2}\big{)}\big{(}1+n^{2}\alpha^{2}x^{2}\big{)}-4n^{2}\alpha\,rx\,.\end{split}$ (36)
When the accelerating parameter is set to zero, the above type-D NUT C-metric apparently reduces to the Taub-NUT solution, and also it obviously recovers the usual C-metric when the NUT charge vanishes. There is a natural question remained, namely, whether it is superficially accelerating just as pointed out in Ref. CQG22-3467 for the traditional form. Then, if it is so, we still lack a type-D NUT C-metric. However, this case seems unlikely. So, we will assume that the above NUT C-metric is not isometric to the Taub-NUT solution in the remaining context, but clearly an obvious proof is still needed.
The above solution has also been generalized in Ref. 2404.06551 to include the electric and magnetic charges as well as a nonzero cosmological constant. However, it suffices enough for our aim to turn off these three parameters. Also, in our opinion, it becomes useful only in the non-rotating case, as we will demonstrate the above rotating and accelerating solution can be transformed into the traditional form of the general P-D metric.
II.5 Griffiths-Podolsky’s parametrization
We next transform from Weyl-Lewis-Papapetrou coordinates ($\rho,z$) to spherical-like coordinates ($\tilde{r},\tilde{x}$) via:
$\rho=\frac{\sqrt{\widetilde{\Delta}(\tilde{r})\widetilde{P}(\tilde{x})}}{[1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})]^{2}}\,,\qquad z=\frac{\big{[}\tilde{x}+\tilde{\alpha}\tilde{r}(\tilde{a}+\tilde{n}\tilde{x})\big{]}\big{[}\tilde{r}-\tilde{m}+\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})\big{(}\tilde{m}\tilde{r}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}\big{]}}{\big{[}1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})\big{]}^{2}}\,,$
where
$\begin{split}\widetilde{\Delta}(\tilde{r})&=\big{(}\tilde{r}^{2}-2\tilde{m}\tilde{r}+\tilde{a}^{2}-\tilde{n}^{2}\big{)}\big{[}(1+\tilde{\alpha}\tilde{n}\tilde{r})^{2}-\tilde{\alpha}^{2}\tilde{a}^{2}\tilde{r}^{2}\big{]}\,,\\ \widetilde{P}(\tilde{x})&=\big{(}1-\tilde{x}^{2}\big{)}\big{[}1+2\tilde{m}\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})+\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\tilde{\alpha}^{2}(\tilde{n}+\tilde{a}\tilde{x})^{2}\big{]}\,,\end{split}$ (37)
in which $\tilde{\sigma}=\sqrt{\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}}$.
The locations of three rod endpoints are now taken to be
$z_{1}=-\tilde{\sigma},\quad z_{2}=\tilde{\sigma},\quad z_{3}=\frac{1+2\tilde{\alpha}\tilde{m}\tilde{n}-\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}{2\tilde{\alpha}\tilde{a}}\,,$
and their corresponding solitons are also purely algebraic:
$\mu_{1}$ $=$ $\frac{[1-\tilde{\alpha}(\tilde{a}-\tilde{n})\tilde{r}](\tilde{r}-\tilde{m}-\tilde{\sigma})[1+(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})](1-\tilde{x})}{[1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})]^{2}}\,,$
$\mu_{2}$ $=$ $\frac{[1-\tilde{\alpha}(\tilde{a}-\tilde{n})\tilde{r}](\tilde{r}-\tilde{m}+\tilde{\sigma})[1+(\tilde{m}+\tilde{\sigma})\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})](1-\tilde{x})}{[1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})]^{2}}\,,$ (38)
$\mu_{3}$ $=$ $\frac{[1+\tilde{\alpha}(\tilde{a}+\tilde{n})\tilde{r}][1-\tilde{\alpha}(\tilde{a}-\tilde{n})\tilde{r}][1+(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})][1+(\tilde{m}+\sigma)\tilde{\alpha}(\tilde{n}+\tilde{a}\tilde{x})]}{\tilde{\alpha}[1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})]^{2}}\,.$
It is troublesome to convert the four functions ($A,B,C,H$) into the expressions in terms of spherical-like coordinates. A simple trick to do this relatively easy is to avoid using the square-root, such as, exploiting identities: $\tilde{\sigma}^{2k}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}$ and $\tilde{\sigma}^{2k+1}=\big{(}\tilde{m}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}^{k}\tilde{\sigma}$ can efficiently facilitate to rationalize the expressions containing the square-root. This experience is also very effective to manipulate the phantom soliton whose square-root function could not be reduced to a simple algebraic expression.
In order to let the conformal factor include a familiar multiplier $\widetilde{\Sigma}=\tilde{r}^{2}+(\tilde{n}+\tilde{a}\tilde{x})^{2}$, we have to take
$C_{1}=\frac{(\tilde{a}-\tilde{n})[1+(\tilde{m}+\tilde{\sigma})\tilde{\alpha}(\tilde{a}+\tilde{n})]}{(\tilde{m}+\tilde{\sigma})[1-(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{a}-\tilde{n})]}\,,\quad C_{2}=\frac{(\tilde{a}+\tilde{n})[1-(\tilde{m}+\tilde{\sigma})\tilde{\alpha}(\tilde{a}-\tilde{n})]}{(\tilde{m}+\tilde{\sigma})[1+(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{a}+\tilde{n})]}\,,$ (39)
and
$k=\frac{(\tilde{m}+\tilde{\sigma})[1+(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{a}+\tilde{n})][1-(\tilde{m}-\tilde{\sigma})\tilde{\alpha}(\tilde{a}-\tilde{n})]}{2(\tilde{\alpha}\tilde{a})^{3/2}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}}\,.$ (40)
Finishing the remaining task of the algebraic process, we find that it is instructive to cast the metric into the Boyer-Lindquist-like form rather than the Lewis-Papapetrou form, which is rather concise and simple enough:
$d\tilde{s}^{2}=\frac{1}{\big{[}1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})\big{]}^{2}}\bigg{\{}\frac{-\widetilde{\Delta}(\tilde{r})\tilde{X}^{2}+\widetilde{P}(\tilde{x})\tilde{Y}^{2}}{\widetilde{\Sigma}}+\widetilde{\Sigma}\Big{[}\frac{d\tilde{r}^{2}}{\widetilde{\Delta}(\tilde{r})}+\frac{d\tilde{x}^{2}}{\widetilde{P}(\tilde{x})}\Big{]}\bigg{\}}\,,$ (41)
where $\widetilde{\Sigma}=\tilde{r}^{2}+(\tilde{n}+\tilde{a}\tilde{x})^{2}$ and
$\begin{split}\tilde{X}&=\frac{\big{[}1-\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\tilde{\alpha}^{2}(\tilde{n}+\tilde{a}\tilde{x})^{2}\big{]}dt+\tilde{\alpha}\tilde{a}\big{[}2\tilde{n}\tilde{x}+\tilde{a}\big{(}1+\tilde{x}^{2}\big{)}\big{]}d\varphi}{\sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}}\,,\\ \tilde{Y}&=\frac{\tilde{a}\big{[}1+\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\tilde{\alpha}^{2}\tilde{r}^{2}\big{]}dt-\tilde{\alpha}\tilde{a}\big{(}\tilde{r}^{2}+\tilde{n}^{2}-\tilde{a}^{2}\big{)}d\varphi}{\sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}}\,.\end{split}$ (42)
In the above frame, the Rindler background or the accelerating horizons is rest relative to the conformal infinity.
One can note that the above metric resembles somewhat like the one derived in the last subsection. In particular, they only differ by an overall constant factor in the case without the NUT charge and by further making identifications: $\alpha=\tilde{\alpha}\tilde{a}$, $m=\tilde{m}$, $a=\tilde{a}$, $r=\tilde{r}$, $x=\tilde{x}$.
The final step that we need to recast the above solution into the familiar form of the general P-D metric is to make the following linear combinations of the Killing coordinates:
$t=\frac{\sqrt{\tilde{\alpha}\tilde{a}}(\tilde{t}-2\tilde{a}\tilde{\phi})}{\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}}\,,\qquad\varphi=\frac{\tilde{a}\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\tilde{t}+\big{[}1-\tilde{\alpha}^{2}\big{(}\tilde{a}^{4}-\tilde{n}^{4}\big{)}\big{]}\tilde{\phi}}{\sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}}\,,$
which yield the familiar expressions:
$\tilde{X}=d\tilde{t}+\big{[}2\tilde{n}\tilde{x}-\tilde{a}\big{(}1-\tilde{x}^{2}\big{)}\big{]}d\tilde{\phi}\,,\quad\tilde{Y}=\tilde{a}\,d\tilde{t}-\big{(}\tilde{r}^{2}+\tilde{a}^{2}+\tilde{n}^{2}\big{)}d\tilde{\phi}\,.$
Note that in this new frame, the Rindler horizon is rotating relative to the conformal infinity.
To summarize, we have re-derived the accelerating Kerr-NUT solution via the ISM procedure from the Rindler background seed. Then using two different parameterizations, we have recast the solution into two distinct forms in terms of spherical-like coordinates, namely, the one novelly given by Ref. PRD109-084038 , and another by the traditional one. Since the latter is already known to be of type-D, so is the one delivered by Astorino in PRD109-084038 . However, one still needs to find what coordinate transformations that may relate them. We leave this subject to the next section.
III Möbius transformations amongst two metrics
In order to find the relation between the above two metrics, a direct routine is to examine the solutions given by Eqs. (33) and (41), respectively, since both of them adopt the same Killing coordinates ($t,\varphi$). Therefore, the remaining most possible coordinate transformation is the Möbius transformations amongst the radial and azimuthal coordinates, respectively, together with further parameter identifications.
Let’s consider the following Möbius transformations:
$\tilde{x}=\frac{x-\nu}{1-\nu\,x}\,,\qquad\tilde{\alpha}\tilde{r}=\frac{\alpha\,r+\nu}{\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)}\,,$ (43)
[…]
$\alpha\,r=\frac{(\tilde{a}-\nu\tilde{n})\tilde{\alpha}\tilde{r}-\nu}{1+(\tilde{n}-\nu\tilde{a})\tilde{\alpha}\tilde{r}}\,.$ (44)
Here and hereafter we prefer to leave $\nu$ to be a freely variable parameter so that the relation between the two mass parameters is left to be determined only in the final step.
We first apply the above transformations to $\widetilde{\Sigma}=\tilde{r}^{2}+(\tilde{n}+\tilde{a}\tilde{x})^{2}$ and find its relation to $\Sigma$. To do so, it is easy to find that if the identity $\widetilde{\Sigma}(1-\nu\,x)^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}=C\,\Sigma$ holds true, then we must have the following conditions:
$\begin{split}&\frac{n}{a}=-\,\frac{\big{(}1+\nu^{2}\big{)}\tilde{n}-2\nu\tilde{a}}{\big{(}1+\nu^{2}\big{)}\tilde{a}-2\nu\tilde{n}}\,,\qquad a^{2}C=\frac{\tilde{a}(\tilde{n}-\nu\tilde{a})(\tilde{a}-\nu\tilde{n})\big{[}\big{(}1+\nu^{2}\big{)}\tilde{a}-2\nu\tilde{n}\big{]}^{2}}{\big{(}1+\nu^{2}\big{)}\tilde{n}-2\nu\tilde{a}}\,,\\ &\tilde{\alpha}^{2}=\frac{-\nu}{(\tilde{n}-\nu\tilde{a})(\tilde{a}-\nu\tilde{n})\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}}\,,\qquad\alpha^{2}a^{2}=\frac{-\nu(\tilde{a}-\nu\tilde{n})\big{[}\big{(}1+\nu^{2}\big{)}\tilde{a}-2\nu\tilde{n}\big{]}^{2}}{\big{(}1-\nu^{2}\big{)}^{2}(\tilde{n}-\nu\tilde{a})\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}}\,.\end{split}$ (45)
Next, we seek the links between the 1-forms ($\tilde{X},\tilde{Y}$) in (42) and their counterparts ($X,Y$) without rescaling the Killing coordinates:
$X$ $=$ $\big{[}1-\big{(}a^{2}-n^{2}\big{)}\alpha^{2}x^{2}\big{]}\frac{dt}{\sqrt{\alpha}}+\big{[}2nx-a\big{(}1+x^{2}\big{)}\big{]}\sqrt{\alpha}d\varphi\,,$
$Y$ $=$ $\big{[}a\big{(}1+\alpha^{2}r^{2}\big{)}+2n\alpha\,r\big{]}\frac{dt}{\sqrt{\alpha}}+\big{(}r^{2}+n^{2}-a^{2}\big{)}\sqrt{\alpha}d\varphi\,.$
It is relatively easy to examine $\sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}$ ($\tilde{X},\tilde{Y}$) rather than ($\tilde{X},\tilde{Y}$). We further find
$\begin{split}\sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}(1-\nu\,x)^{2}\tilde{X}&=\frac{\tilde{a}\sqrt{\alpha}\big{(}1-\nu^{2}\big{)}}{\tilde{a}-\nu\tilde{n}}X\,,\\ \sqrt{\tilde{\alpha}\tilde{a}}\sqrt{1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}\tilde{Y}&=-\frac{\tilde{a}^{2}\alpha^{3/2}\big{(}1-\nu^{2}\big{)}}{\tilde{\alpha}(\tilde{a}-\nu\tilde{n})}Y\,,\end{split}$ (46)
and must additionally impose
$a=-\,\frac{\tilde{\alpha}(\tilde{a}-\nu\tilde{n})\big{[}\big{(}1+\nu^{2}\big{)}\tilde{a}-2\nu\tilde{n}\big{]}}{\alpha\big{(}1-\nu^{2}\big{)}}\,,$ (47)
then we can write
$n=\frac{\tilde{\alpha}(\tilde{a}-\nu\tilde{n})\big{[}\big{(}1+\nu^{2}\big{)}\tilde{n}-2\nu\tilde{a}\big{]}}{\alpha\big{(}1-\nu^{2}\big{)}}\,,\quad C=-\alpha^{2}\frac{\tilde{a}(\tilde{n}-\nu\tilde{a})^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\big{(}1-\nu^{2}\big{)}^{2}}{\nu\big{[}\big{(}1+\nu^{2}\big{)}\tilde{n}-2\nu\tilde{a}\big{]}}\,.$ (48)
In practice, if we first find the links between ($\tilde{X},\tilde{Y}$) and ($X,Y$) directly, then all the above relations can be obtained, and it can be checked that the identity $\widetilde{\Sigma}(1-\nu\,x)^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}=C\,\Sigma$ is also automatically fulfilled.
Note that the relation for $\alpha^{2}$ listed above then becomes the identity for $\tilde{\alpha}^{2}$, so $\alpha$ can now temporarily be viewed as a free parameter too. Alternatively, we can also express this in term of untilded parameters as follows:
$\alpha^{2}=\frac{\nu(a-\nu\,n)}{(n-\nu\,a)\big{(}a^{2}-n^{2}\big{)}}\,.$
Our next step is to find the relation between the mass parameters. Using the following identities:
$1-\tilde{x}^{2}=\frac{1-\nu^{2}}{(1-\nu\,x)^{2}}\big{(}1-x^{2}\big{)}\,,\quad\frac{d\tilde{x}^{2}}{1-\tilde{x}^{2}}=\frac{1-\nu^{2}}{(1-\nu\,x)^{2}}\,\frac{dx^{2}}{1-x^{2}}\,,$
we can simply let
$\frac{d\tilde{x}^{2}}{\widetilde{P}(\tilde{x})}=\big{(}1-\nu^{2}\big{)}\frac{dx^{2}}{P(x)}\,,$ (49)
so that after using the above relation for $\tilde{\alpha}^{2}$, we find that this is true only when we further set
$\tilde{m}=\frac{\nu}{2\tilde{\alpha}(\tilde{a}-\nu\tilde{n})}\,,\qquad m=\frac{\nu\big{(}\nu^{2}\tilde{a}-3\tilde{a}+2\nu\tilde{n}\big{)}}{2\alpha(\tilde{a}-\nu\tilde{n})}\,,$ (50)
which can also be attained by exploiting the radial part.
Finally, with the help of the following useful expressions:
$1+\tilde{\alpha}\tilde{r}(\tilde{n}+\tilde{a}\tilde{x})=\frac{\tilde{a}\big{(}1-\nu^{2}\big{)}(1+\alpha\,rx)}{(1-\nu\,x)[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]}\,,\quad\widetilde{\Sigma}=\frac{C\,\Sigma}{(1-\nu\,x)^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{2}}\,,$
$\widetilde{\Delta}(\tilde{r})=\frac{\big{(}1-\nu^{2}\big{)}\tilde{a}^{2}\alpha^{2}}{\tilde{\alpha}^{2}[\tilde{a}(1+\nu\alpha\,r)-\tilde{n}(\alpha\,r+\nu)]^{4}}\Delta(r)\,,\quad\widetilde{P}(\tilde{x})=\frac{1-\nu^{2}}{(1-\nu\,x)^{4}}P(x)\,,\quad\frac{d\tilde{r}^{2}}{\widetilde{\Delta}(\tilde{r})}=\big{(}1-\nu^{2}\big{)}\frac{dr^{2}}{\Delta(r)}\,,$
and the relations (46) between ($\tilde{X},\tilde{Y}$) and ($X,Y$), we can show that the two metrics only differ by an overall constant factor:
$d\tilde{s}^{2}$ $=$ $\frac{C\,\Sigma}{\big{(}1-\nu^{2}\big{)}\tilde{a}^{2}(1+\alpha\,rx)^{2}}\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}+\frac{\tilde{a}\alpha^{3}\big{(}1-\nu^{2}\big{)}}{C\tilde{\alpha}^{3}(\tilde{a}-\nu\tilde{n})^{2}\big{[}1+\tilde{\alpha}^{2}\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}^{2}]}\,\frac{-\Delta(r)X^{2}+P(x)Y^{2}}{(1+\alpha\,rx)^{2}\Sigma}$ (51)
$=$ $\frac{C}{\tilde{a}^{2}\big{(}1-\nu^{2}\big{)}}\bigg{\{}\frac{\Sigma}{(1+\alpha\,rx)^{2}}\Big{[}\frac{dr^{2}}{\Delta(r)}+\frac{dx^{2}}{P(x)}\Big{]}+\frac{n}{(1-\nu^{2})(\tilde{n}-\nu\tilde{a})}\frac{-\Delta(r)X^{2}+P(x)Y^{2}}{(1+\alpha\,rx)^{2}\Sigma}\bigg{\}}$
$=$ $\frac{C\big{(}1+\alpha^{2}a^{2}\big{)}}{\tilde{a}^{2}\big{(}1-\nu^{2}\big{)}}\,ds^{2}\,,$
provided that we set
$1+\alpha^{2}a^{2}=\frac{n}{(1-\nu^{2})(\tilde{n}-\nu\tilde{a})}$
which yields an equation to determine $\alpha$:
$\frac{\tilde{\alpha}}{\alpha}\big{[}\big{(}1+\nu^{2}\big{)}\tilde{n}-2\nu\tilde{a}\big{]}=\big{(}1-\nu^{2}\big{)}^{2}\frac{\tilde{n}-\nu\tilde{a}}{\tilde{a}-\nu\tilde{n}}-\frac{\nu}{\tilde{a}^{2}-\tilde{n}^{2}}[\big{(}1+\nu^{2}\big{)}\tilde{a}-2\nu\tilde{n}\big{]}^{2}\,.$ (52)
Thus we have generally proven that the two metrics given in the last two sections are isometric to each other. It seems likely to let $C=\tilde{a}^{2}\big{(}1-\nu^{2}\big{)}/\big{(}1+\alpha^{2}a^{2}\big{)}$ so that both metrics are completely identical. This condition implies
$\frac{\tilde{\alpha}^{2}}{\alpha^{2}}=\frac{\big{(}\nu^{4}+1\big{)}\tilde{a}^{2}-\nu\big{(}3\nu^{2}+1\big{)}\tilde{a}\tilde{n}+\big{(}3\nu^{2}-1\big{)}\tilde{n}^{2}}{\tilde{a}(\tilde{a}-\nu\tilde{n})\big{(}\tilde{a}^{2}-\tilde{n}^{2}\big{)}\big{(}1-\nu^{2}\big{)}}\,,$ (53)
when combined with Eq. (52), they yield a cubic equation for $\tilde{a}/\tilde{n}$:
$\nu(\nu^{4}-\nu^{2}+1)\tilde{a}^{3}-\nu^{2}(3\nu^{2}-2)\tilde{a}^{2}\tilde{n}+\nu(3\nu^{2}-4)\tilde{a}\tilde{n}^{2}+\tilde{n}^{3}=0\,.$ (54)
One can also conduct a similar task as done in the above by using the C-metric-like coordinates ($x,y$) with $y=-1/(\alpha\,r)$ and ($\bar{x},\bar{y}$) defined via $\bar{y}=-1/(\tilde{\alpha}\tilde{r})$, $\bar{x}=\tilde{n}+\tilde{a}\tilde{x}$, rather than by employing the spherical-like coordinates. In this case, both metrics become more symmetric, and now $\tilde{a}$ behaves like a scale parameter, for example, $\bar{y}=\tilde{n}+\tilde{a}(y-\nu)/(1-\nu\,y)$. Since the results remain the same as those given in the above, we will not repeat this work although it might be relatively easy.
What limit should be taken in the original P-D metric to get the type-D NUT C-metric?
Previously, Griffiths and Podolsky CQG22-3467 demonstrated that in the nonrotating limit, the NUT C-metric is apparently accelerating and can be transformed into the Taub-NUT solution, as a result of this, the type-D NUT C-metric seems to be absent from the most general family of the type-D P-D solution, leaving a puzzle for almost two decades. The reason for this is probably that they didn’t consider the Möbius transformation. On the other hand, it is a simple matter to just set $a=0$ in the Astorino’s expression PRD109-084038 to arrive at the type-D C-metric, but this still does not explain how to get the type-D NUT C-metric from the most general P-D solution.
To educate the lesson why the type-D NUT C-metric is missing from the P-D metric, let’s consider what limit we should take in the original P-D metric so as to get the expected type-D NUT C-metric. Let’s first assume that the transformation parameter $\nu$ and all the solution parameters ($\tilde{m},\tilde{n},\tilde{a},\tilde{\alpha}$) as well as the radial coordinate $\tilde{r}$ are all finite and nonzero (although $\tilde{a}$ can be set to zero). To achieve the non-rotating NUT C-metric, we must first let the rotation parameter $a$ tend to zero ($a\to 0$). Then from the first identity in (45), the NUT charge parameter $n$ must also tend to a very tiny constant ($n\simeq 0$) so that $n/a$ keeps finite. From the fourth one in (45) or (47), we know that the acceleration parameter $\alpha$ must approach the infinity ($\alpha\to\infty$) so that $\alpha\,a$ is also finite. The second identity in (45) tells us that the overall factor $C$ should tend to the infinity ($C\to\infty$) so that $C\alpha$ can be kept finite. From the inverse Möbius transformation (44), we can recognize that the radial coordinate $r$ must approach zero so that $\alpha\,r$ remains finite. Finally, from the second Eq. (50), we learn that the mass parameter $m$ should tend to a very tiny constant $(m\simeq 0)$. In the $a\to 0$ limit, both ($\alpha\,a$, $\alpha\,n$, $\alpha\,m$) and $\alpha\,r$ remains fixed.
To summarize, in order to get the type-D NUT C-metric, the corresponding limit is the following set:
$a\to 0\,,\quad n\simeq 0\,,\quad m\simeq 0\,,\quad r\to 0\,,\qquad\alpha\to\infty\,,$ (55)
apart from the appropriate rescaling of the whole line element plus the linear combinations of the Killing coordinates ($t\,,\varphi$). However, the key measure is to consider the above Möbius transformations (43).
IV Transforming the type-D NUT C-metric into the familiar P-D form
In this section, we will consider another obvious coordinate transformations that bring the type-D NUT C-metric (35) into the familiar P-D form. To this end, let’s introduce the following coordinate transformations:
$t=4n\mu(\tilde{t}+\tilde{\phi})\,,\quad\varphi=4n\mu\alpha(\tilde{t}-\tilde{\phi})\,,\quad r=n\frac{q+1}{q-1}\,,\quad x=\frac{1-p}{n\alpha(1+p)}\,,\quad m=\frac{n}{2\mu}\,,$ (56)
which bring the NUT C-metric (35) into the P-D form:
$ds^{2}=\frac{8n^{2}\mu}{(p-q)^{2}}\bigg{\{}\frac{Q(q)\big{(}d\tilde{t}+p^{2}d\tilde{\phi}\big{)}^{2}+P(p)\big{(}q^{2}d\tilde{t}-d\tilde{\phi}\big{)}^{2}}{1+p^{2}q^{2}}+\big{(}1+p^{2}q^{2}\big{)}\Big{[}\frac{dq^{2}}{-Q(q)}+\frac{dp^{2}}{P(p)}\Big{]}\bigg{\}}\,,$ (57)
where the structure function reads
$P(\xi)=\big{[}n^{2}\alpha^{2}(1+\xi)^{2}-(1-\xi)^{2}\big{]}\big{(}1+4\mu\xi-\xi^{2}\big{)}\,,$ (58)
in which $\alpha$ can be simply set to unity. The only nonzero Kretschmann invariant is:
$R^{abcd}R_{abcd}$ $=$ $\frac{3\alpha^{4}(p-q)^{6}}{\mu^{2}\tilde{\nu}^{4}(1+p^{2}q^{2})^{6}}\big{[}2\mu\big{(}1-\tilde{\nu}^{2}\big{)}pq\big{(}p^{2}q^{2}-3\big{)}+\big{(}\tilde{\nu}^{2}+1\big{)}\big{(}3p^{2}q^{2}-1\big{)}\big{]}$ (59)
×$\qquad\times\big{[}2\mu\big{(}1-\tilde{\nu}^{2}\big{)}\big{(}3p^{2}q^{2}-1\big{)}-\big{(}\tilde{\nu}^{2}+1\big{)}pq\big{(}p^{2}q^{2}-3\big{)}\big{]}\,,$
where $\tilde{\nu}=n\alpha$. This Kretschmann invariant has a similar factorized structure form as that of the Taub-NUT solution, but it does not coincide with the latter because the former contains a conformal factor $(p-q)^{6}$. It demonstrates that the type-D NUT C-metric can not reduce to the Taub-NUT solution unless the acceleration parameter vanishes ($\alpha=0$).
From the above expression for the structure function, it is clear that this quartic function is factorized in an entirely distinct manner from the traditional way. This perhaps is the primary reason why the type-D NUT C-mtric is “lost” from the most general P-D solution.
V Concluding remarks
In this paper, we have shown that the type-D accelerating Kerr-NUT metric presented in Ref. PRD109-084038 is not a new solution, but it is still very useful in the non-rotating case, namely, the type-D static accelerating solution with a nonzero NUT charge since this solution is unavailable in the previous literature. We have justified this point from two aspects. First, we have applied the ISM procedure to generate the accelerating Kerr-NUT solution from the seed metric – Rindler vacuum background. Then we adopted two different parameterizations to derive the metric given by Astorino in Ref. PRD109-084038 and the traditional one early delivered by Griffiths and Podolsky. This clearly demonstrates that these two metrics are isometric, though displayed in two different forms. Second, we have also provided the obvious coordinate transformations (Möbius transformations of the radial and angular coordinates and linear combinations of the Killing coordinates) and parameter identifications between these two line elements. In particular, we also presented another concrete Möbius transformation and the linear combinations of the Killing coordinates that obviously cast the type-D NUT C-metric into the familiar form of the P-D solution. We anticipate that this work can give a final answer to resolve the “missing” puzzle of the type-D NUT C-metric.
In the appendix Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme, we have given two simple examples to illustrate our new normalization scheme by deriving the 5-dimensional Schwarzschild-Tangherlini solution from the flat vacuum via the ISM construction. Our normalization strategy can be viewed as a methodological complementary to the Pomeransky’s “remove–re-add” tricky.
Note added: When this paper is under review (while arXiv on-holds our submission on August 27), an eprint 2409.02308 appears with a partial overlap of the subject for which the Möbius transformation is also used to relate different forms of the accelerating Kerr-NUT solution. The Möbius transformation previously used in Ref. PRD93-044021 was suggested to Dr. M. Astorino by S.-Q. Wu in an email communication about five months ago.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12375053, No. 12205243, No. 11675130, by the Sichuan Science and Technology Program under Grant No. 2023NSFSC1347, and by the Doctoral Research Initiation Project of China West Normal University under Grant No. 21E028.
Appendix: ISM construction of 5-dimensional static solutions with our new normalization scheme
In this appendix, we will show that by using our normalization scheme, the 5-dimensional Schwarzschild-Tangherlini solution can be generated from the flat vacuum background via the ISM construction, contrary to the statement made in Ref. PRD73-044004 . Our derivation strictly follows the spirt of the ISM construction, and is different from that had been done in Ref. PTP114-793 where the Killing metric is modified.
V.1 Generating 5-dimensional Schwarzschild-Tangherlini solution
We first take the 5-dimensional flat Minkowski background
$ds^{2}=-dt^{2}+\rho^{2}d\phi^{2}+d\psi^{2}+d\rho^{2}+dz^{2}\,,$
as our seed metric:
$\mathbf{g}_{0}=\mathop{\rm diag}\nolimits(-1\,,\rho^{2}\,,1)\,,$ (A.1)
so the generating matrix is easily written as
$\mathbf{\Psi}_{0}=\mathop{\rm diag}\nolimits(-1\,,\rho^{2}-2\lambda\,z-\lambda^{2}\,,1)\,.$
[…]
$\mathbf{m}=\bigg{(}\begin{array}[]{ccc}1&0&0\\ 0&\frac{1}{\mu_{2}}&0\end{array}\bigg{)}\,,\qquad\mathbf{l}=\bigg{(}\begin{array}[]{ccc}-1&0&0\\ 0&\frac{\rho^{2}}{\mu_{2}}&0\end{array}\bigg{)}\,,\qquad\mathbf{\Gamma}=\bigg{(}\begin{array}[]{cc}\frac{-1}{\rho^{2}+\mu_{1}^{2}}&0\\ 0&\frac{\rho^{2}}{\mu_{2}^{2}\big{(}\rho^{2}+\mu_{1}^{2}\big{)}}\end{array}\bigg{)}\,.$ (A.8)
The generated 5-dimensional static 2-soliton metric is very simple:
$\widetilde{\mathbf{g}}=\mathop{\rm diag}\nolimits\Big{(}\frac{\rho^{2}}{\mu_{1}^{2}}\,,\frac{-\rho^{4}}{\mu_{2}^{2}}\,,1\Big{)}\,,$ (A.9)
with the determinant $\det\widetilde{\mathbf{g}}=-\rho^{6}/\big{(}\mu_{1}^{2}\mu_{2}^{2}\big{)}\not=-\rho^{2}$.
It is suggested to take the normalization diagonal matrix as
$\mathbf{W}=\mathop{\rm diag}\nolimits\Big{(}\frac{\mu_{1}^{3/2}}{\rho\sqrt{\mu_{2}}}\,,\frac{\mu_{2}}{\rho\sqrt{\mu_{1}}}\,,\sqrt{-\mu_{2}}\Big{)}\,,$ (A.10)
then the final physics metric is exactly the expected one:
$\mathbf{g}=-\mathbf{W}\widetilde{\mathbf{g}}\mathbf{W}=-\mathbf{W}^{2}\widetilde{\mathbf{g}}=\mathop{\rm diag}\nolimits\Big{(}\frac{-\mu_{1}}{\mu_{2}}\,,\frac{\rho^{2}}{\mu_{1}}\,,\mu_{2}\Big{)}\,,$ (A.11)
with the conformal factor being recorded as
$f=\frac{\mu_{2}\big{(}\rho^{2}+\mu_{1}\mu_{2}\big{)}}{\big{(}\rho^{2}+\mu_{1}^{2}\big{)}\big{(}\rho^{2}+\mu_{2}^{2}\big{)}}\,.$ (A.12)
This trivial example illuminates that the 5-dimensional Schwarzschild-Tangherlini solution is feasibly generated via ISM procedure with our new normalization scheme from the flat Minkowski vacuum seed.
V.2 Generating 5-dimensional static Emparan-Reall black ring
Next, we want to re-derive the static Emparan-Reall black ring PRL88-101101 solution from the 5-dimensional flat Minkowski background written in another form:
$ds^{2}=-dt^{2}+\frac{\rho^{2}}{\mu_{3}}d\phi^{2}+\mu_{3}d\psi^{2}+\frac{\mu_{3}}{\rho^{2}+\mu_{3}^{2}}\big{(}d\rho^{2}+dz^{2}\big{)}\,.$ (B.1)
Clearly, its $t={\rm\,const}$ slice is the Euclidean Rindler metric.
From the seed metric
$\mathbf{g}_{0}=\mathop{\rm diag}\nolimits\Big{(}-1\,,\frac{\rho^{2}}{\mu_{3}}\,,\mu_{3}\Big{)}\,,$ (B.2)
we rewrite $\rho^{2}/\mu_{3}=-\nu_{3}$ and make the replacements: $\mu_{3}\to\mu_{3}-\lambda$, $\nu_{3}\to\nu_{3}-\lambda$, then easily obtain the generating matrix:
$\mathbf{\Psi}_{0}=\mathop{\rm diag}\nolimits\Big{(}-1\,,\frac{\rho^{2}}{\mu_{3}}+\lambda\,,\mu_{3}-\lambda\Big{)}\,.$
[…]
$\mathbf{m}=\bigg{(}\begin{array}[]{ccc}-1&0&0\\ 0&\frac{\mu_{3}}{\rho^{2}+\mu_{2}\mu_{3}}&0\end{array}\bigg{)}\,,\quad\mathbf{l}=\bigg{(}\begin{array}[]{ccc}1&0&0\\ 0&\frac{\rho^{2}}{\rho^{2}+\mu_{2}\mu_{3}}&0\end{array}\bigg{)}\,,\quad\mathbf{\Gamma}=\bigg{(}\begin{array}[]{cc}\frac{-1}{\rho^{2}+\mu_{1}^{2}}&0\\ 0&\frac{\rho^{2}\mu_{3}}{\big{(}\rho^{2}+\mu_{2}^{2}\big{)}\big{(}\rho^{2}+\mu_{2}\mu_{3}\big{)}^{2}}\end{array}\bigg{)}\,.$
The generated 5-dimensional 2-soliton metric is still diagonal:
$\widetilde{\mathbf{g}}=\mathop{\rm diag}\nolimits\Big{(}\frac{\rho^{2}}{\mu_{1}^{2}}\,,\frac{-\rho^{4}}{\mu_{2}^{2}\mu_{3}}\,,\mu_{3}\Big{)}\,,$ (B.3)
with the determinant $\det\widetilde{\mathbf{g}}=-\rho^{6}/\big{(}\mu_{1}^{2}\mu_{2}^{2}\big{)}\not=-\rho^{2}$.
Now we suggest to take the normalization matrix as
$\mathbf{W}=\mathop{\rm diag}\nolimits\Big{(}\frac{\mu_{1}^{3/2}}{\rho\sqrt{\mu_{2}}}\,,\frac{\mu_{2}^{3/2}}{\rho\sqrt{\mu_{1}}}\,,\sqrt{-1}\Big{)}\,,$ (B.4)
then the final physics metric is exactly that of the static Emparan-Reall black ring:
$\mathbf{g}=-\mathbf{W}\widetilde{\mathbf{g}}\mathbf{W}=-\mathbf{W}^{2}\widetilde{\mathbf{g}}=\mathop{\rm diag}\nolimits\Big{(}\frac{-\mu_{1}}{\mu_{2}}\,,\frac{\rho^{2}\mu_{2}}{\mu_{1}\mu_{3}}\,,\mu_{3}\Big{)}\,,$ (B.5)
whose conformal factor can be easily computed as
$f=\frac{k\,\mu_{3}\big{(}\rho^{2}+\mu_{1}\mu_{2}\big{)}^{2}\big{(}\rho^{2}+\mu_{2}\mu_{3}\big{)}}{(1-\nu)\big{(}\rho^{2}+\mu_{1}\mu_{3}\big{)}\big{(}\rho^{2}+\mu_{1}^{2}\big{)}\big{(}\rho^{2}+\mu_{2}^{2}\big{)}\big{(}\rho^{2}+\mu_{3}^{2}\big{)}}\,.$ (B.6)
In the coalescing limit $\mu_{3}\mapsto\mu_{2}$ with $k=1-\nu$, the solution reduces to the 5-dimensional Schwarzschild-Tangherlini solution. On the other hand, in the merging limit $\mu_{2}\mapsto\mu_{1}$ with $k=1-\nu$, the solution coalesces to the above 5-dimensional Minkowski background metric.